a dog is 21ft from the door. it runs toward the door at a constant velocity. after 5s, it is 1 from the door. what was the dogs velocity?


1
Expert's answer
2021-10-06T08:14:09-0400
"v=\dfrac{s}{t}=\dfrac{21\ ft-1\ ft}{5\ s}=4\ \dfrac{ft}{s}."